The equipotential surfaces in the field map belowThe equipotential surfaces in the field map below (shown as dashed lines) have potential differences between a point at infinity and themselves of 1V, 5V, and 10V. (a)[4 pt(s) ]Label the surfaces with the correct voltage. Justify your choices using the properties of electric potential and its relation to work. (b)[3 pt(s) ]Is the work to move a +q charge from points A to B positive, negative, or about zero? Justify using the properties of work and electric potential. (c)[3 pt(s) ]Is the work to move a +q charge from points A to C positive, negative, or about zero? Justify using the properties of work and electric potential. DRAW THE PATH TO WHICH YOUR REASONING APPLIES.
Answer #1

a)
Potential can be written as V = kQ/r
It depends both on the charge and the distance from the charge. Here in all the cases the net charge is +Q. So the potential depends on the distance from the charge only.
10 V corresponds to the innermost equipotential surface since r is minimum
5 V corresponds to the equipotential surface in the middle
1 V corresponds to the outermost equipotential surface

b)
Work done = q[VB - VA]
Since VA and VB have the same potential,
VA - VB = 5 - 5 = 0
Work done = 0.

c)
Work done = q[VC - VA]
q[1 - 5]
= -4q
Work done is negative.